# Dream-in-Style: Text-to-3D Generation using Stylized Score Distillation

### v1

@@ Abstract.
Text-to-3D Generation, 3D Stylization, Diffusion Models, Score Distillation
ccs: Computing methodologies Shape modelingccs: Computing methodologies Shape inferenceccs: Computing methodologies Image-based rendering
Figure 1. We aim to generate a 3D object jointly from a text prompt and a style reference image so that the object integrates the descriptive elements in the text prompt and the aesthetic style in the reference image. Our method employs a stylized score distillation to steer a text-to-3D optimization using combined scores from a pretrained text-to-image model and its modified variant on attention layer features, generating stylized 3D objects in a single-stage optimization.
1. Introduction
Creating 3D content has been a key but demanding task in computer graphics. Traditional interactive tools such as Maya (Autodesk, 2019), Blender (Blender Foundation, 2018) are among the most popular choices for novices and professionals to perform 3D modeling. In the wave of generative AI development, there have been increased interests in automatic synthesis of 3D content using generative models (Poole et al., 2023; Liu et al., 2023). This is an open research area with tremendous progress in recent years, with the rise of language models enabling the widespread adoption of natural languages to condition the automatic generation of data in different modalities.
This trend has stimulated the development of text-to-3D generation methods (Poole et al., 2023; Wang et al., 2023a), where 3D objects can be generated by simply prompting an input sentence that describes the desired object content. These methods are generic to the appearance of the generated objects, which means that the final look and feel of the 3D content is barely controllable. This is in contrast to a common requirement in traditional 3D modeling, where a visual artist might aim to decorate a 3D object in particular styles. For example, one might be interested in creating a 3D object with low polygon count, making its geometry appear as a collection of flat surfaces, or a 3D object with stylized textures in photorealistic or cartoon styles. Performing such a stylization using traditional tools is a tedious task. Therefore, integrating stylization into generative models is a promising idea to explore.
At its core, text-to-3D generation (Poole et al., 2023; Wang et al., 2023b) performs an iterative update on a 3D representation such that its rendering converges to a photorealistic image scored by a pretrained text-to-image model. So far, most development of text-to-3D generation has focused on objects with generic appearance. Creating 3D content with particular styles remains challenging to achieve. A straightforward approach is to incorporate style description into text prompts used to generate 3D content, but this approach is not effective due to the ambiguity in how styles can be described using natural languages.
@@ 1. Introduction
In summary, our contributions are threefold. First, we propose to adapt a generic pretrained text-to-image model with a reference style image and build a training-free modified pretrained model for stylized text-to-3D generation. Second, we propose a stylized score distillation gradient to steer the 3D generation toward the desired style specified in the reference style image. Third, we demonstrate the flexibility of our method by applying our stylized distillation to different text-to-3D generation losses, including score distillation sampling (SDS) (Poole et al., 2023), noise-free score distillation (Katzir et al., 2024), and variational score distillation (VSD) (Wang et al., 2023b). We also demonstrate the robustness of our 3D stylization via a diverse set of text prompts and styles.
2. Related works
2.1. Style transfer
@@
Style transfer is a traditional computer graphics problem that aims to synthesize images in artistic styles. Example-based methods such as image analogies (Hertzmann et al., 2001; Liao et al., 2017) learn to filter a pair of source images so that when applied to a new target image, the filter can generate analogous filtered results. Image analogies are built upon texture synthesis, which requires source pairs to approximately align to learn effective filters. This restriction is alleviated in modern deep-learning based style transfer methods. Neural style transfer (Gatys et al., 2016) optimizes the target image so that it shares the content of an input image and the style of a reference image with feature correspondences characterized by a pretrained neural network. Several extensions of style transfer follow, notably for efficiency improvement with a feed-forward neural network (Johnson et al., 2016) and image-to-image translation networks (Isola et al., 2017; Zhu et al., 2017; Liu et al., 2017), style representation with statistical features (Huang and Belongie, 2017), style representation using text-image features (Gal et al., 2022; Patashnik et al., 2021; Kwon and Ye, 2022), and video style transfer (Li et al., 2019). We refer the reader to the comprehensive survey paper (Jing et al., 2019) for a broader coverage on visual style transfer methods.
2.2. Generative content creation
@@
Generative models such as generative adversarial networks (Goodfellow et al., 2014; Karras et al., 2019) and diffusion models (Ho et al., 2020; Po et al., 2024) are notable tools to synthesize realistic data of different modalities, notably text and images when trained on large-scale datasets. Recently, text-to-image models like DALL-E (OpenAI, 2023a) and Stable Diffusion (Rombach et al., 2022) have shown great promise in generating photorealistic images from arbitrary text prompts. Text-to-image diffusion models can be used for image editing (Meng et al., 2022) including style transfer by learning a copy of a pretrained diffusion model (Zhang et al., 2023), bridging the latent space of two diffusion models for image-to-image translations (Su et al., 2023; Li et al., 2023c), and instruction-based translations supervised using image-prompt-image datasets (Brooks et al., 2023) and test-time editing directions (Parmar et al., 2023). Personalized text-to-image methods (Ruiz et al., 2023) can generate images in similar subjects defined by a small set of reference images, but these methods requires multiple images to finetune the pretrained diffusion models. Textual inversion methods (Gal et al., 2023; Han et al., 2023) instead only optimize the text embedding to obtain a text prompt that preserves the subject identity of an input image. Training-free methods (Hertz et al., 2024; Jeong et al., 2024) inject reference features into the denoising process of a pretrained diffusion model through manipulating features input to attention layers of the denoising U-Net to influence the stylized generation.
@@
In the 3D domain, generative models can be trained to sample 3D data represented by voxels (Wu et al., 2015) and point clouds (Nichol et al., 2022), but high-quality 3D training datasets are relatively scarce and are in smaller scales compared to language and image datasets (Deitke et al., 2023b, a). A recent approach to sidestep this issue is to generate 3D data by learning from images. 3D-aware GANs (Chan et al., 2022) learn to generate 3D-consistent images by incorporating a neural radiance field (Mildenhall et al., 2020) as the intermediate 3D representation, but their results are limited to a few categories of objects. Instead of focusing on 3D generation, novel view synthesis (Chan et al., 2023; Szymanowicz et al., 2023) predicts 3D-consistent views from a sparse set of input views, with support from pretrained image diffusion models (Liu et al., 2023). Large reconstruction models (Hong et al., 2024; Xu et al., 2024) demonstrate effective image-to-3D generation by scaling up the training to millions of 3D objects. Text-to-3D generation (Poole et al., 2023; Wang et al., 2023a, b) is a recent advance aiming to generate 3D objects directly from text prompts by using a pretrained text-to-image model to score the rendering of 3D objects at random angles. Text-to-3D methods have witnessed rapid development recently, with significant advances made toward improved distillation (Katzir et al., 2024), shape quality (Lin et al., 2023; Xu et al., 2023; Chen et al., 2023a; Zhu et al., 2024) with textures (Metzer et al., 2023), fast rendering (Tang et al., 2024), amortized sampling (Lorraine et al., 2023; Xie et al., 2024), 3D editing (Hertz et al., 2023; Koo et al., 2024), and animated models (Bahmani et al., 2024). Our method belongs to the text-to-3D family, but focuses on stylized 3D generation.
2.3. 3D stylization
Several methods for 3D stylization exist. Traditionally, image analogies can be adapted to stylize 3D rendering while preserving physically based illumination effects (Fišer et al., 2016; Sýkora et al., 2019). These methods assume that the target 3D object is available, and only the texture and illumination on the target are changed during style transfer.
By using neural radiance fields (NeRFs) (Mildenhall et al., 2020), a straightforward method to integrate style transfer with 3D content creation is to adapt a style loss and a content loss during NeRF reconstruction using an alternating optimization strategy (Nguyen-Phuoc et al., 2022). This method is extended to improved stylization quality by optimizing semantic correspondences between the radiance field and the style reference image (Zhang et al., 2022; Pang et al., 2023), assuming a two-stage process involving a NeRF reconstruction and then stylization. In this work, we aim to integrate styles into 3D generation by considering style guidance using reference images in a text-to-3D generation framework. Our method differs from NeRF stylization in that we focus on single-stage generation that simultaneously optimizes 3D geometry, appearance, and styles. We refer the reader to a recent survey (Chen et al., 2023b) for further details on 3D stylization methods.
3. Background
3.1. Text-to-3D generation
@@
The basic concept in text-to-3D generation is to use a pretrained text-to-image diffusion model to score the rendering of a 3D object described by a text prompt. Particularly, given a text prompt $y$ and a pretrained text-to-image diffusion model with the noise prediction network $\epsilon_{\phi}(\mathbf{z}_{t}\mid y)$, we aim to generate a 3D object, parameterized by $\theta$, such that its rendering $\mathbf{x}=g(\theta)$ follows the image distribution generated by the pretrained diffusion model. This generation can be formulated as an optimization problem with a score distillation sampling (SDS) gradient (Poole et al., 2023):
@@
(1) $\nabla_{\theta}\mathcal{L}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[\omega(t)\left(\epsilon_{\phi}(\mathbf{z}_{t}\mid y)-\epsilon\right)\frac{\partial\mathbf{x}}{\partial\theta}\right],$
@@
where $\omega(t)$ is a weighting function with $t\sim\mathcal{U}(0.02,0.098)$, $\epsilon\sim\mathcal{N}(0,I)$, $\mathbf{z}_{t}=\alpha_{t}\mathbf{x}+\sigma_{t}\epsilon$. In practice, the score function $\epsilon_{\phi}(\mathbf{z}_{t}\mid y)$ is implemented with classifier-free guidance (Ho and Salimans, 2021) to steer the denoising process toward conditional generation to align the generated samples with text prompt $y$. Several variants of score distillation have been explored to improve the fidelity of generated 3D objects, e.g., variational score distillation (VSD) that expresses the generated objects as probabilistic distributions (Wang et al., 2023b), noise-free score distillation (Katzir et al., 2024) that decomposes distillation scores into interpretable components.
3.2. Baseline methods for 3D stylization
@@
Neural style loss
The challenges encountered from the first method motivate us to use a reference image to describe detailed visual elements for style transfer. Our second baseline method involves using a neural style loss (Gatys et al., 2016) to enforce style consistency between the 3D rendering and a style reference image. The style loss is defined by
(2) $\mathcal{L}_{style}(\theta)=\|f(\mathbf{x})-f(\mathbf{s})\|_{2}^{2},$
where $f$ represents the style features extracted by VGG-19 (Simonyan and Zisserman, 2015). Style features were extracted from images using the conv1_1, conv2_1, conv3_1, conv4_1, conv4_2, and conv5_1 layers. We apply the style loss as a regularization to an existing score distillation loss.
Textual inversion
The third baseline method is specialized for text-to-3D generation by using textual inversion (Gal et al., 2023; Han et al., 2023) to map a style reference image to the text embedding of a text-to-image pretrained model, resulting in an augmented text prompt that implicitly encodes the style reference image. Particularly, we follow (Gal et al., 2023) to optimize a token $h$ to reproduce the style reference image $\mathbf{s}$ so that the augmented prompt can be defined by $y^{\prime}=[y+\text{``in the style of''}+h]$. We can then use the augmented prompt $y^{\prime}$ instead of $y$ in a standard text-to-3D optimization. This baseline method depends on the accuracy of textual inversion that might affect the final 3D generation, and also requires additional computation to perform the textual inversion.
@@
4. Method
4.1. Overview
Our method seeks a 3D object such that its rendering aligns to an input text prompt and a style reference image. We optimize a 3D neural representation using score distillation, where the rendering of the 3D object is scored by a pretrained text-to-image diffusion model (Poole et al., 2023). Compared to the generic text-to-3D generation, one particular challenge here is to integrate the style reference image into the optimization process to generate stylized 3D objects. Our method is designed to be a single-stage optimization, where both stylized geometry and appearance are generated simultaneously. This differs from some existing 3D stylization methods where only geometry or appearance is optimized to stylize a pre-constructed neural representation (Zhang et al., 2022; Pang et al., 2023).
We propose to consume our style reference image using an attention swapping mechanism on the denoising U-Net of the pretrained diffusion model (Jeong et al., 2024; Hertz et al., 2024) so that the modified diffusion model can generate images analogously to the style reference image. We show that this modified pretrained model remains suitable for score distillation, which we then leverage to guide the 3D optimization.
4.2. Style-based score distillation
@@
(3) $\mathcal{L}(\theta)=KL(q(\mathbf{z}_{t}\mid\mathbf{x}=g(\theta))\parallel p_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s})),$
@@
To model $p_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s})$, we take inspiration from training-free methods for style transfer using diffusion models (Hertz et al., 2024; Jeong et al., 2024). We assume that there are two denoising processes: one process for generating an image using the original text prompt, and another process for generating a style reference image. Here, the style reference image can be generated by its own text prompt, or from textual inversion of a real style image. Our goal is to influence the former process so that its generated image has the original content but shares the style in the latter process. This can be achieved by sharing features in self-attention blocks (Hertz et al., 2024) or swapping key and value features at self-attention blocks of the latter process with those of the original process (Jeong et al., 2024), allowing features from the style reference images to propagate into image synthesis of the former process. As no finetuning is done on the diffusion model itself, this leaves the parameters of the original diffusion model intact, only the score predictions are updated due to the feature changes in the self-attention blocks. We adopt this concept for text-to-3D generation as it allows us to use the same pretrained model for original and stylized score distillation for text-to-3D generation. In our implementation, we follow the swapped attention in visual style prompting (Jeong et al., 2024) but similar methods such as shared attention (Hertz et al., 2024) should work as well.
@@
(4) $\hat{\epsilon}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s})=\epsilon_{\phi}(\mathbf{z}_{t}\mid y;att(y_{s})),$
@@
(5) $\log p_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s})=(1-\lambda)\log p_{\phi}(\mathbf{z}_{t}\mid y)+\lambda\log\hat{p}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s}),$
@@
(6) $\left.-(1-\lambda)\log(p_{\phi}(\mathbf{z}_{t}\mid y))-\lambda\log(\hat{p}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s}))\right].$
@@
(7) $\nabla_{\theta}\mathcal{L}_{SSD}$ $=\mathbb{E}_{t,\epsilon}\left[\omega(t)\left((1-\lambda)\epsilon_{\phi}(\mathbf{z}_{t}\mid y)+\lambda\hat{\epsilon}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s})-\epsilon\right)\frac{\partial\mathbf{x}}{\partial\theta}\right].$
Notably, our stylized score distillation results in linearly interpolated scores of the original and modified pretrained diffusion model, resembling SDS-family gradients. This makes extensions on SDS become applicable on our method as well, e.g., classifier-free guidance (Ho and Salimans, 2021), and noise-free score distillation (Katzir et al., 2024), as demonstrated subsequently.
@@
Following the noise-free score distillation loss decomposition and taking into account classifier-free guidance (Katzir et al., 2024), we can represent the score function as a composition of a domain direction $\delta_{D}$, a noise direction $\delta_{N}$, and a conditioning direction $\delta_{C}$. The noise-free version of our stylized score distillation can be written as
(8) $\nabla_{\theta}\mathcal{L}_{SNF}=\mathbb{E}_{t}\left[\omega(t)\left((1-\lambda)\left(\delta_{D}+\beta\delta_{C}\right)+\lambda\left(\hat{\delta}_{D}+\beta\hat{\delta}_{C}\right)\right)\frac{\partial\mathbf{x}}{\partial\theta}\right],$
@@
(9) $\delta_{D}=\begin{cases}\epsilon_{\phi}(\mathbf{z}_{t}\mid y=\emptyset),&\text{if}~{}~{}t<200\\ \epsilon_{\phi}(\mathbf{z}_{t}\mid y=\emptyset)-\epsilon_{\phi}(\mathbf{z}_{t}\mid y=p_{neg}),&\text{otherwise},\\ \end{cases}$
@@
(10) $\hat{\delta}_{D}=\begin{cases}\hat{\epsilon}_{\phi}(\mathbf{z}_{t}\mid y=\emptyset,\mathbf{s}),&\text{if}~{}~{}t<200\\ \hat{\epsilon}_{\phi}(\mathbf{z}_{t}\mid y=\emptyset,\mathbf{s})-\hat{\epsilon}_{\phi}(\mathbf{z}_{t}\mid y=p_{neg},\mathbf{s}),&\text{otherwise},\\ \end{cases}$
@@
(11) $\delta_{C}=\epsilon_{\phi}(\mathbf{z}_{t}\mid y)-\epsilon_{\phi}(\mathbf{z}_{t}\mid y=\emptyset),$
@@
(12) $\hat{\delta}_{C}=\hat{\epsilon}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s})-\hat{\epsilon}_{\phi}(\mathbf{z}_{t}\mid y=\emptyset,\mathbf{s}),$

### v2

@@ Abstract
1 Introduction
Creating 3D content has been a key but demanding task in computer graphics. Traditional interactive tools such as Maya [1], Blender [3] are among the most popular choices for novices and professionals to perform 3D modeling. In the wave of generative AI development, there have been increased interests in automatic synthesis of 3D content using generative models [59, 45]. This is an open research area with tremendous progress in recent years, with the rise of language models enabling the widespread adoption of natural languages to condition the automatic generation of data in different modalities.
This trend has stimulated the development of text-to-3D generation methods [59, 70], where 3D objects can be generated by simply prompting an input sentence that describes the desired object content. These methods are generic to the appearance of the generated objects, which means that the final look and feel of the 3D content is barely controllable. This is in contrast to a common requirement in traditional 3D modeling, where a visual artist might aim to decorate a 3D object in particular styles. For example, one might be interested in creating a 3D object with low polygon count, making its geometry appear as a collection of flat surfaces, or a 3D object with stylized textures in photorealistic or cartoon styles. Performing such a stylization using traditional tools is a tedious task. Therefore, integrating stylization into generative models is a promising idea to explore.
At its core, text-to-3D generation [59, 72] performs an iterative update on a 3D representation such that its rendering converges to a photorealistic image scored by a pretrained text-to-image model. So far, most development of text-to-3D generation has focused on objects with generic appearance. Creating 3D content with particular styles remains challenging to achieve. A straightforward approach is to incorporate style description into text prompts used to generate 3D content, but this approach is not effective due to the ambiguity in how styles can be described using natural languages.
Figure 1: We aim to generate a 3D object jointly from a text prompt and a style reference image so that the object integrates the descriptive elements in the text prompt and the aesthetic style in the reference image. Our method employs a stylized score distillation to steer a text-to-3D optimization using combined scores from a pretrained text-to-image model and its modified variant on attention layer features, generating stylized 3D objects in a single-stage optimization.
@@ 1 Introduction
In summary, our contributions are threefold. First, we propose to adapt a generic pretrained text-to-image model with a reference style image and build a training-free modified pretrained model for stylized text-to-3D generation. Second, we propose a stylized score distillation gradient to steer the 3D generation toward the desired style specified in the reference style image. Third, we demonstrate the flexibility of our method by applying our stylized distillation to different text-to-3D generation losses, including score distillation sampling (SDS) [59], noise-free score distillation [34], and variational score distillation (VSD) [72]. We also demonstrate the robustness of our 3D stylization via a diverse set of text prompts and styles.
2 Related works
2.1 Style transfer
@@
Style transfer is a traditional computer graphics problem that aims to synthesize images in artistic styles. Example-based methods such as image analogies [23, 41] learn to filter a pair of source images so that when applied to a new target image, the filter can generate analogous filtered results. Image analogies are built upon texture synthesis, which requires source pairs to approximately align to learn effective filters. This restriction is alleviated in modern deep-learning based style transfer methods. Neural style transfer [17] optimizes the target image so that it shares the content of an input image and the style of a reference image with feature correspondences characterized by a pretrained neural network. Several extensions of style transfer follow, notably for efficiency improvement with a feed-forward neural network [32] and image-to-image translation networks [29, 81, 44], style representation with statistical features [27], style representation using text-image features [15, 57, 36], and video style transfer [40]. We refer the reader to the comprehensive survey paper [31] for a broader coverage on visual style transfer methods.
2.2 Generative content creation
@@
Generative models such as generative adversarial networks [18, 33] and diffusion models [25, 58] are notable tools to synthesize realistic data of different modalities, notably text and images when trained on large-scale datasets. Recently, text-to-image models like DALL-E [53] and Stable Diffusion [61] have shown great promise in generating photorealistic images from arbitrary text prompts. Text-to-image diffusion models can be used for image editing [48] including style transfer by learning a copy of a pretrained diffusion model [80], bridging the latent space of two diffusion models for image-to-image translations [65, 37], and instruction-based translations supervised using image-prompt-image datasets [4] and test-time editing directions [56]. Personalized text-to-image methods [62] can generate images in similar subjects defined by a small set of reference images, but these methods requires multiple images to finetune the pretrained diffusion models. Textual inversion methods [16, 20] instead only optimize the text embedding to obtain a text prompt that preserves the subject identity of an input image. Training-free methods [22, 30] inject reference features into the denoising process of a pretrained diffusion model through manipulating features input to attention layers of the denoising U-Net to influence the stylized generation.
@@
In the 3D domain, generative models can be trained to sample 3D data represented by voxels [75] and point clouds [52], but high-quality 3D training datasets are relatively scarce and are in smaller scales compared to language and image datasets [11, 10]. A recent approach to sidestep this issue is to generate 3D data by learning from images. 3D-aware GANs [5] learn to generate 3D-consistent images by incorporating a neural radiance field [50] as the intermediate 3D representation, but their results are limited to a few categories of objects. Instead of focusing on 3D generation, novel view synthesis [6, 67] predicts 3D-consistent views from a sparse set of input views, with support from pretrained image diffusion models [45]. Large reconstruction models [26, 78] demonstrate effective image-to-3D generation by scaling up the training to millions of 3D objects. Text-to-3D generation [59, 70, 72] is a recent advance aiming to generate 3D objects directly from text prompts by using a pretrained text-to-image model to score the rendering of 3D objects at random angles. Text-to-3D methods have witnessed rapid development recently, with significant advances made toward improved distillation [34], shape quality [42, 77, 7, 82] with textures [49], fast rendering [68], amortized sampling [47, 76], 3D editing [21, 35], and animated models [2]. Our method belongs to the text-to-3D family, but focuses on stylized 3D generation.
2.3 3D stylization
Several approaches for 3D stylization exist. Traditionally, image analogies can be adapted to stylize 3D rendering while preserving physically based illumination effects [14, 66]. Recent notable advances in neural radiance fields (NeRFs) [50] allow us to perform stylization on implicit neural scene representations. NeRF stylization methods [51, 13, 43, 79, 55] often assume a two-stage process in which a radiance field is first reconstructed on photorealistic images and then stylized based on style reference images. This process can be improved by several techniques including optimizing semantic correspondences between the radiance field and the style reference image [79, 55] or generalizing the stylization across scenes and styles with generalizable NeRFs [9, 28]. Two-stage stylization methods tend to change only the appearance of the NeRF while keeping the geometry intact [9, 28] because these methods lack a generative prior and therefore cannot generate style-related geometry. For example, given a car model and a fire style image in Fig. 1, two-stage NeRF stylization methods cannot generate a car made of fire with additional geometry representing fire. We refer the reader to a recent survey [8] for more techniques on 3D stylization.
With the rapid development of image-to-3D based on large reconstruction models [26, 73, 69, 46, 78], one can also consider generating stylized 3D objects by lifting a single stylized image to 3D. However, these pretrained image-to-3D models do not generalize well to stylized images with complex visual effects, e.g., fire around a car, resulting in 3D models with limited geometry quality. By contrast, our method integrates styles into 3D generation by considering style guidance using reference images in a text-to-3D generation framework. Our method follows a single-stage generation principle, simultaneously optimizing 3D geometry, appearance, and styles.
3 Background
3.1 Text-to-3D generation
@@
The basic concept in text-to-3D generation is to use a pretrained text-to-image diffusion model to score the rendering of a 3D object described by a text prompt. Particularly, given a text prompt $y$ and a pretrained text-to-image diffusion model with the noise prediction network $\epsilon_{\phi}(\mathbf{z}_{t}\mid y)$, we aim to generate a 3D object, parameterized by $\theta$, such that its rendering $\mathbf{x}=g(\theta)$ follows the image distribution generated by the pretrained diffusion model. This generation can be formulated as an optimization problem with a score distillation sampling (SDS) gradient [59]:
@@
$\nabla_{\theta}\mathcal{L}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[\omega(t)\left(\epsilon_{\phi}(\mathbf{z}_{t}\mid y)-\epsilon\right)\frac{\partial\mathbf{x}}{\partial\theta}\right],$ (1)
@@
where $\omega(t)$ is a weighting function with $t\sim\mathcal{U}(0.02,0.098)$, $\epsilon\sim\mathcal{N}(0,I)$, $\mathbf{z}_{t}=\alpha_{t}\mathbf{x}+\sigma_{t}\epsilon$. In practice, the score function $\epsilon_{\phi}(\mathbf{z}_{t}\mid y)$ is implemented with classifier-free guidance [24] to steer the denoising process toward conditional generation to align the generated samples with text prompt $y$. Several variants of score distillation have been explored to improve the fidelity of generated 3D objects, e.g., variational score distillation (VSD) that expresses the generated objects as probabilistic distributions [72], noise-free score distillation [34] that decomposes distillation scores into interpretable components.
3.2 Baseline methods for 3D stylization
@@
Neural style loss.
The challenges encountered from the first method motivate us to use a reference image to describe detailed visual elements for style transfer. Our second baseline method involves using a neural style loss [17] to enforce style consistency between the 3D rendering and a style reference image. The style loss is defined by
$\mathcal{L}_{style}(\theta)=\|f(\mathbf{x})-f(\mathbf{s})\|_{2}^{2},$ (2)
where $f$ represents the style features extracted by VGG-19 [64]. Style features were extracted from images using the conv1_1, conv2_1, conv3_1, conv4_1, conv4_2, and conv5_1 layers. We apply the style loss as a regularization to an existing score distillation loss.
Textual inversion.
The third baseline method is specialized for text-to-3D generation by using textual inversion [16, 20] to map a style reference image to the text embedding of a text-to-image pretrained model, resulting in an augmented text prompt that implicitly encodes the style reference image. Particularly, we follow [16] to optimize a token $h$ to reproduce the style reference image $\mathbf{s}$ so that the augmented prompt can be defined by $y^{\prime}=[y+\text{``in the style of''}+h]$. We can then use the augmented prompt $y^{\prime}$ instead of $y$ in a standard text-to-3D optimization. This baseline method depends on the accuracy of textual inversion that might affect the final 3D generation, and also requires additional computation to perform the textual inversion.
@@
4 Method
4.1 Overview
Our method seeks a 3D object such that its rendering aligns to an input text prompt and a style reference image. We optimize a 3D neural representation using score distillation, where the rendering of the 3D object is scored by a pretrained text-to-image diffusion model [59]. Compared to the generic text-to-3D generation, one particular challenge here is to integrate the style reference image into the optimization process to generate stylized 3D objects. Our method is designed to be a single-stage optimization, where both stylized geometry and appearance are generated simultaneously. This differs from some existing 3D stylization methods where only geometry or appearance is optimized to stylize a pre-constructed neural representation [79, 55].
We propose to consume our style reference image using an attention swapping mechanism on the denoising U-Net of the pretrained diffusion model [30, 22] so that the modified diffusion model can generate images analogously to the style reference image. We show that this modified pretrained model remains suitable for score distillation, which we then leverage to guide the 3D optimization.
4.2 Style-based score distillation
@@
$\mathcal{L}(\theta)=KL(q(\mathbf{z}_{t}\mid\mathbf{x}=g(\theta))\parallel p_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s})),$ (3)
@@
To model $p_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s})$, we take inspiration from training-free methods for style transfer using diffusion models [22, 30]. We assume that there are two denoising processes: one process for generating an image using the original text prompt, and another process for generating a style reference image. Here, the style reference image can be generated by its own text prompt, or from textual inversion of a real style image. Our goal is to influence the former process so that its generated image has the original content but shares the style in the latter process. This can be achieved by sharing features in self-attention blocks [22] or swapping key and value features at self-attention blocks of the latter process with those of the original process [30], allowing features from the style reference images to propagate into image synthesis of the former process. As no finetuning is done on the diffusion model itself, this leaves the parameters of the original diffusion model intact, only the score predictions are updated due to the feature changes in the self-attention blocks. We adopt this concept for text-to-3D generation as it allows us to use the same pretrained model for original and stylized score distillation for text-to-3D generation. In our implementation, we follow the swapped attention in visual style prompting [30] but similar methods such as shared attention [22] should work as well.
@@
$\hat{\epsilon}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s})=\epsilon_{\phi}(\mathbf{z}_{t}\mid y;att(y_{s})),$ (4)
@@
$\log p_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s})=(1-\lambda)\log p_{\phi}(\mathbf{z}_{t}\mid y)+\lambda\log\hat{p}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s}),$ (5)
@@
$\left.-(1-\lambda)\log(p_{\phi}(\mathbf{z}_{t}\mid y))-\lambda\log(\hat{p}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s}))\right].$ (6)
@@
$\nabla_{\theta}\mathcal{L}_{SSD}$ $=\mathbb{E}_{t,\epsilon}\left[\omega(t)\left((1-\lambda)\epsilon_{\phi}(\mathbf{z}_{t}\mid y)\right.\right.$
$\left.\left.+\lambda\hat{\epsilon}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s})-\epsilon\right)\frac{\partial\mathbf{x}}{\partial\theta}\right].$ (7)
Notably, our stylized score distillation results in linearly interpolated scores of the original and modified pretrained diffusion model, resembling SDS-family gradients. This makes extensions on SDS become applicable on our method as well, e.g., classifier-free guidance [24], and noise-free score distillation [34], as demonstrated subsequently.
@@
Following the noise-free score distillation loss decomposition and taking into account classifier-free guidance [34], we can represent the score function as a composition of a domain direction $\delta_{D}$, a noise direction $\delta_{N}$, and a conditioning direction $\delta_{C}$. The noise-free version of our stylized score distillation can be written as
$\nabla_{\theta}\mathcal{L}_{SNF}$ $=\mathbb{E}_{t}\left[\omega(t)\left((1-\lambda)\left(\delta_{D}+\beta\delta_{C}\right)\right.\right.$
$\left.\left.+\lambda\left(\hat{\delta}_{D}+\beta\hat{\delta}_{C}\right)\right)\frac{\partial\mathbf{x}}{\partial\theta}\right],$ (8)
@@
$\delta_{D}=\begin{cases}\epsilon_{\phi}(\mathbf{z}_{t}\mid y=\emptyset),&\text{if}~{}~{}t<200\\ \epsilon_{\phi}(\mathbf{z}_{t}\mid y=\emptyset)-\epsilon_{\phi}(\mathbf{z}_{t}\mid y=p_{neg}),&\text{otherwise},\\ \end{cases}$ (9)
@@
$\hat{\delta}_{D}=\begin{cases}\hat{\epsilon}_{\phi}(\mathbf{z}_{t}\mid y=\emptyset,\mathbf{s}),&\text{if}~{}~{}t<200\\ \hat{\epsilon}_{\phi}(\mathbf{z}_{t}\mid y=\emptyset,\mathbf{s})-\hat{\epsilon}_{\phi}(\mathbf{z}_{t}\mid y=p_{neg},\mathbf{s}),&\text{otherwise},\\ \end{cases}$ (10)
@@
$\delta_{C}=\epsilon_{\phi}(\mathbf{z}_{t}\mid y)-\epsilon_{\phi}(\mathbf{z}_{t}\mid y=\emptyset),$ (11)
@@
$\hat{\delta}_{C}=\hat{\epsilon}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s})-\hat{\epsilon}_{\phi}(\mathbf{z}_{t}\mid y=\emptyset,\mathbf{s}),$ (12)